# TREASUREHUNT: Transients and Variability Discovered with HST in the JWST North Ecliptic Pole Time Domain Field

### v2

@@ 4.1.1 Transients with Significant X-ray Detections
For the XMM-Newton and NuSTAR catalogs, we used a matching radius of 5″ and 30″, respectively, and found a likely match for transients T7, T10, and T11. We inspected the Chandra 0.5–7 keV event-file images for obvious sources, and here also identified sources consistent with T7 and T10. For the other transients, we extracted source counts from a region with a radius of 2″ and background counts from a nearby source-free region with a radius of 10″, but found no evidence for any X-ray counterparts above 1 significance. Chandra upper limits vary across the field due to detector geometry, as well as vignetting and PSF spreading for off-axis sources, but we expect typical 3 upper limits of 1.410 ct s for the full 1.3 Ms exposure. For T7 and T10, we then extracted source fluxes using the ciao tool srcflux on the cross-registered event files for each OBSID. To convert count rates to fluxes, we used a model with assumed Galactic absorption and a = 1.4 power law. At 6.5′ from the field center (characteristic of T7 and T10), this corresponds to 3 upper limits of 3.210 erg s cm. Note, however, that the epochs of individual X-ray observations do not necessarily match or overlap with specific HST visits in which a transient was observed or absent.
@@
[erg s cm] | [keV] | ||
T7 | Chandra | (5.360.64)10 | 0.5–7 |
T7 | XMM-Newton | (1.410.41)10 | 0.5–2 |
T7 | XMM-Newton | 9.03 10 | 2–10 |
T10 | Chandra | (2.680.53)10 | 0.5–7 |
T10 | XMM-Newton | (1.260.42)10 | 0.5–2 |
T10 | XMM-Newton | 4.97 10 | 2–10 |
T11 | NuSTAR | (1.140.41)10 | 3–8 |
@@ 4.2 Variability in the JWST NEP Time Domain Field
We also estimate the areal density of variable sources using the total number of inferred genuine variables. These 110 unique variable candidates translate to 1.25 variables per arcmin (4500 deg) of overlapping area. We can therefore expect 14 new variable sources per additional fully overlapping ACS/WFC footprint (11.3 arcmin), assuming F606W imaging to similar depths as the TREASUREHUNT observations.
@@ 4.3 Photometric Redshift Estimation and SED Fits
We created photometric catalogs for input into EAZY using SourceExtractor. Recognizing that some galaxies exhibit variability that may bias a redshift estimation, we ran SourceExtractor on HST drizzled images of each individual visit instead of the full HST mosaic, and use the measurement with the fainter flux. We assume the fainter measurement represents the photometry if no variability were present —galaxies may temporarily brighten, but they cannot get dimmer on human time scales. For the JWST data, we used the full mosaics, since only a single epoch yet exists at any given location within the field. The NIRCam F444W image served as the SourceExtractor detection image in all cases. For HST images, we set MAG_ZEROPOINT to 26.49 mag for F606W and to 25.65 mag for F435W. For our JWST images, MAG_ZEROPOINT was set to 28.0865 mag for all filters, as appropriate for their 0030 platescale, to convert from MJy sr (see Windhorst et al., 2023). Other relevant parameters are the same for HST and JWST, and are as listed in the right-most column of Table 1.
@@
Due to the necessarily limited grid of SED model parameters, any galaxy with an inferred sSFR 10 yr is categorized as “quiescent” and its sSFR is set to 10 yr. Our templates also cannot distinguish stellar population ages younger than 10 Myr, so we set all apparent age solutions 0.01 Gyr to 0.01 Gyr. Given that age is inversely correlated with the inferred sSFR, where Age corresponds to sSFR, we also set all (sSFR) to 2. This adjustment is solely made to assess collective trends in galaxies with variability.
In Figure 8, we compare photometric redshift (), stellar mass, stellar population age, specific star formation rate (sSFR), and dust extinction () of galaxies that exhibit variability to typical galaxies in the general population. Overall, we sample a broad range of redshifts, masses, ages, star formation histories, and dust extinction. The most prominent peaks in the fraction of galaxies with variability with respect to the general galaxy population (expressed in %) occur at a redshift of 2.5, a stellar mass of 10 and 10 , an age of 10 Gyr, a sSFR of 10 Gyr, and an attenuation 1.0 mag. The median redshift of our variability sample is = 1.3, but we sample variability up to = 2.9 . Similarly, our galaxies with variability sample masses = (0.02–20)10 , ages = (0.01–2) Gyr, sSFR = (10–10) Gyr, and = (0.0–2.4) mag.
@@
The best fit SED for the T4 host galaxy is shown in Figure 9 and corresponds to a stellar population age of 3.6 Gyr, = 0.5 mag of attenuation by dust, and a stellar mass of 1.210 . Its sSFR is 0.5 Gyr. The reported best-fit value of 0.05 of the reduced indicates that the (conservative) photometric uncertainties upon input to CIGALE were overestimating the actual ones somewhat.
@@ 5.1 Classification of Transients
The TREASUREHUNT observing strategy provided piecemeal UV–Visible imaging rather than time-domain monitoring. As a result, these SNe candidates will have long faded and can no longer be followed up for spectroscopic identification. Nonetheless, now that this initial imaging exists, future monitoring observations could efficiently expand the present sample and would allow such spectroscopic follow-up. The sum of the areas of overlap used in this work samples 88 arcmin, which is about half of the total HST coverage of this field (194 arcmin). Therefore, we can expect about 2 as many transients if this field were observed again. Full overlapping coverage of this field could yield at least six new Type Ia supernovae, and coupled with spectra and rapid follow-up, can provide essential constraints on fundamental cosmological parameters like the Hubble constant, the mass density of the universe, the cosmological constant, the deceleration parameter, and the age of the Universe (e.g., Riess et al., 1998, 2023). In addition to Type Ia SNe, comprehensive HST coverage of this field could yield at least six CC SNe observations. Particularly when combined with ancillary data, this could contribute significantly to resolving the observed discrepancies in CC SNe rates compared to theoretical models (Cappellaro, 2014). With a sufficiently large sample, the SNe rate as a function of redshift could also be constrained. Furthermore, the increased volume of observations enhances the likelihood of capturing rare or unexpected events, potentially leading to groundbreaking discoveries in the field.
@@
At least three of the transients are unlikely to be SNe. They either were detected in both epochs or they are isolated with no apparent host galaxy. When considering only the optical emission, transients T7 and T10 exhibit substantial changes in brightness between the two epochs, of = 0.83 0.06 mag and 0.82 0.09 mag, respectively. That would be equivalent to the presence of an additional source with 24.82 and 25.52 mag. The observed centroid positions in both epochs match to within 0001, such that the additional signal is indistinguishable from being due to a nuclear source that changed brightness. We therefore suggest that T7 and T10 may be quasars. The fact that they both were detected in soft X-rays at flux levels 210 erg s cm lends additional evidence to this interpretation.
@@
T11 is a unique transient with no visible host, yet was detected in hard X-rays by NuSTAR. If that X-ray emission is indeed associated with T11 then it could be a faint ( 29 mag), high-redshift quasar that briefly flared up to be detectable in F606W. If the X-ray emission is unrelated, then we speculate that T11 may represent a SN located within an undetected host galaxy. That host could be a faint dwarf galaxy at 6, or a more massive host at 6 to explain the non-detection in F606W in the second visit. Unfortunately, T11 falls outside the area with JWST coverage, so we cannot distinguish between these two scenarios at this time. Chakrabarti et al. (2018) have demonstrated that the rates of SNe in dwarf galaxies can be three times higher (per unit volume) than in typical spiral galaxies. Given the limited understanding of star formation in dwarf galaxies, a substantial number of SNe candidates like T11 could offer a unique opportunity to identify otherwise undetected dwarf galaxies and gain more insights into their star formation rates.
@@ 5.2 Properties of Variable Candidates
The observed stellar mass distribution (Figure 8b) of galaxies exhibiting variability shows an apparent excess at both 10 and 10 . We do not detect variability in the most massive galaxies, either due to small number statistics (most likely) or because their SMBHs are less likely to accrete at a sufficient rate to be detectable or have a longer period of relative quiescence between major accretion events. We also do not detect variability in galaxies with 10 , likely due to a combination of both small number statistics and larger photometric uncertainties in these faint systems, such that our 3 threshold is less likely to be met for a given amplitude of variability. If the relative distribution is taken at face values, then it could also hint at the presence of two separate populations of sources with variability: one associated with lower mass 10 hosts, and another associated with higher mass hosts 10 . Such could be the case if variability associated with CC SNe favors lower mass host galaxies, and variability associated with SN Ia and AGN favors higher mass hosts.
In Figure 8c and d we see that a significant fraction of the galaxies with variability are best fit with stellar population models characterized by active star formation (young ages 30 Myr and high sSFR 30 Gyr). Conversely, there is also a sizeable fraction that is quiescent (sSFR 0.01 Gyr) with stellar populations older than 1 Gyr. A possible excess appears for intermediate population ages (60–600 Myr).
@@ 5.3 Constraints on SMBH Mass, Accretion and Radiation Lifetimes
The sparse and random time sampling of this field does not allow us to trace the evolution of faint AGN, but may still teach us various properties of them. AGN brightness varies on timescales corresponding to the light crossing time in the accretion disks and gas clouds surrounding SMBHs. For example, for a SMBH with a mass 10 , the timescale for variability cannot be shorter than 1 week in the rest-frame (e.g., Xie et al., 2005). For the majority of sources where the variability is due to AGN activity, that variability originates close to the central SMBH. If the timescale we sample is the minimum timescale of variation, we can provide rough constraints on the SMBH masses following Xie et al. (2005). According to Schwarzschild Black Hole Theory (SBHT), the mass of a black hole ( in ) is equal to the minimum timescale of variation ( in yr) as = 4.2910 . Time intervals between observations in various areas of overlap in the TREASUREHUNT data range from 1 day (0.0027 yr) to 4.78 yr. At the median redshift of our variable candidate sample, , the range in rest-frame timescales becomes 0.001–2.08 yr, giving SMBH masses of 5.110 8.910 . SMBH masses 10 are unreasonable: objects within the region of overlap with a 4.78 yr interval between observations likely vary on much shorter timescales and we did not sample the minimum timescale of variation. Even a more typical time interval of 1 yr between epochs will yield a 10 SMBH mass. Taking the full range in redshifts sampled into account, we will therefore only claim that in select areas of overlap we can probe SMBHs with masses of (3–9)10 , but with no firm upper limit.
@@
We can also frame our finding that a maximum of 0.42% of the general field galaxy population shows significant variability in the context of the accretion and radiation lifetimes of AGN. If we assume that: 1) all galaxies seen in this field with HST have a central SMBH (e.g., Kormendy & Ho, 2013), 2) the central SMBH is always visible when accreting, and 3) that a currently accreting AGN will be variable, then we would deduce that the AGN in our field are actively accreting for 0.42% of the time. Assumptions 2) and 3) are simplified, as some AGN will be hidden at rest-frame UV–Visible wavelengths by surrounding dust. This fraction could be as large as 2/3 of all AGN, if we use the 30% average LyC escape fraction of weak AGN (Smith et al., 2018, 2020, 2024) as proxy for the fraction of AGN with direct unobscured sight lines to the observer. Figure 8c implies an average SED age of 10 years (100 Myr) for the parent population, although the spread in age is wide (0.01–10 Gyr). A 0.42% fraction of visibly variable AGN would correspond to an implied average AGN activity lifetime of 0.42%10 410 yr (and a range of 410–410 yr). For comparison, Rawes et al. (2015) estimated the optical synchrotron electron life times of AGN with visible jets observed with HST and Chandra at 10 yr. However, these authors state that their estimated synchrotron lifetimes may be too short by at least a factor of two, and may be longer if synchrotron electrons are re-accelerated in the ambient magnetic field. On the other end of the activity scale, Jakobsen et al. (2003) show that at least some quasars can remain active more or less continuously for 10 Myr. With these assumptions and significant uncertainties, a visible AGN accretion time of 10 to 10 yr could indeed result in a 0.42% variability fraction in a galaxy population characterized by an average SED age of 100 Myr (and a range of 0.01–10 Gyr) whose stars and gas feed that central engine.
@@ 5.4 Caveats and Estimated Population Size
We also note that the variable candidates presented here are a strict lower bound to the number present in the field: many genuine variable sources, even if at 6 and detectable in F606W, were certainly missed because they did not vary sufficiently between specific individual visits to meet our 3 threshold, or if the timescale of their variability exceeded the duration of the TREASUREHUNT program. In the present study, the photometric uncertainties in each epoch of observation contributed equally to that 3 threshold. Any future additional observations to similar depths would reduce the photometric uncertainty in the reference magnitudes, thus lowering the variability detection threshold. The magnitude and level of variability for all sources in this work is collected in Table 5, so that it can be compared in future studies within the JWST NEP TDF.

### v3

@@ 4.1.1 Transients with Significant X-ray Detections
For the XMM-Newton and NuSTAR catalogs, we used a matching radius of 5″ and 30″, respectively, and found a likely match for transients T7, T10, and T11. We inspected the Chandra 0.5–7 keV event-file images for obvious sources, and here also identified sources consistent with T7 and T10. For the other transients, we extracted source counts from a region with a radius of 2″ and background counts from a nearby source-free region with a radius of 10″, but found no evidence for any X-ray counterparts above 1 significance. Chandra upper limits vary across the field due to detector geometry, as well as vignetting and PSF spreading for off-axis sources, but we expect typical 3 upper limits of 1.410-5 ct s-1 for the full 1.3 Ms exposure. For T7 and T10, we then extracted source fluxes using the ciao tool srcflux on the cross-registered event files for each OBSID. To convert count rates to fluxes, we used a model with assumed Galactic absorption and a = 1.4 power law. At 6.5′ from the field center (characteristic of T7 and T10), this corresponds to 3 upper limits of 3.210-16 erg s-1 cm-2. Note, however, that the epochs of individual X-ray observations do not necessarily match or overlap with specific HST visits in which a transient was observed or absent.
@@
[erg s-1 cm-2] | [keV] | ||
T7 | Chandra | (5.360.64)10-15 | 0.5–7 |
T7 | XMM-Newton | (1.410.41)10-15 | 0.5–2 |
T7 | XMM-Newton | 9.03 10-15 | 2–10 |
T10 | Chandra | (2.680.53)10-15 | 0.5–7 |
T10 | XMM-Newton | (1.260.42)10-15 | 0.5–2 |
T10 | XMM-Newton | 4.97 10-15 | 2–10 |
T11 | NuSTAR | (1.140.41)10-14 | 3–8 |
@@ 4.2 Variability in the JWST NEP Time Domain Field
We also estimate the areal density of variable sources using the total number of inferred genuine variables. These 110 unique variable candidates translate to 1.25 variables per arcmin2 (4500 deg-2) of overlapping area. We can therefore expect 14 new variable sources per additional fully overlapping ACS/WFC footprint (11.3 arcmin2), assuming F606W imaging to similar depths as the TREASUREHUNT observations.
@@ 4.3 Photometric Redshift Estimation and SED Fits
We created photometric catalogs for input into EAZY using SourceExtractor. Recognizing that some galaxies exhibit variability that may bias a redshift estimation, we ran SourceExtractor on HST drizzled images of each individual visit instead of the full HST mosaic, and use the measurement with the fainter flux. We assume the fainter measurement represents the photometry if no variability were present —galaxies may temporarily brighten, but they cannot get dimmer on human time scales. For the JWST data, we used the full mosaics, since only a single epoch yet exists at any given location within the field. The NIRCam F444W image served as the SourceExtractor detection image in all cases. For HST images, we set MAG_ZEROPOINT to 26.49 mag for F606W and to 25.65 mag for F435W. For our JWST images, MAG_ZEROPOINT was set to 28.0865 mag for all filters, as appropriate for their 0030 platescale, to convert from MJy sr-1 (see Windhorst et al., 2023). Other relevant parameters are the same for HST and JWST, and are as listed in the right-most column of Table 1.
@@
Due to the necessarily limited grid of SED model parameters, any galaxy with an inferred sSFR 10-5 yr-1 is categorized as “quiescent” and its sSFR is set to 10-5 yr-1. Our templates also cannot distinguish stellar population ages younger than 10 Myr, so we set all apparent age solutions 0.01 Gyr to 0.01 Gyr. Given that age is inversely correlated with the inferred sSFR, where Age corresponds to sSFR, we also set all (sSFR) to 2. This adjustment is solely made to assess collective trends in galaxies with variability.
In Figure 8, we compare photometric redshift (), stellar mass, stellar population age, specific star formation rate (sSFR), and dust extinction () of galaxies that exhibit variability to typical galaxies in the general population. Overall, we sample a broad range of redshifts, masses, ages, star formation histories, and dust extinction. The most prominent peaks in the fraction of galaxies with variability with respect to the general galaxy population (expressed in %) occur at a redshift of 2.5, a stellar mass of 108 and 1010 , an age of 10-0.5 Gyr, a sSFR of 10-5 Gyr-1, and an attenuation 1.0 mag. The median redshift of our variability sample is = 1.3, but we sample variability up to = 2.9 . Similarly, our galaxies with variability sample masses = (0.02–20)109 , ages = (0.01–2) Gyr, sSFR = (10-5–102) Gyr-1, and = (0.0–2.4) mag.
@@
The best fit SED for the T4 host galaxy is shown in Figure 9 and corresponds to a stellar population age of 3.6 Gyr, = 0.5 mag of attenuation by dust, and a stellar mass of 1.21010 . Its sSFR is 0.5 Gyr-1. The reported best-fit value of 0.05 of the reduced indicates that the (conservative) photometric uncertainties upon input to CIGALE were overestimating the actual ones somewhat.
@@ 5.1 Classification of Transients
The TREASUREHUNT observing strategy provided piecemeal UV–Visible imaging rather than time-domain monitoring. As a result, these SN candidates will have long faded and can no longer be followed up for spectroscopic identification. Nonetheless, now that this initial imaging exists, future monitoring observations could efficiently expand the present sample and would allow such spectroscopic follow-up. The sum of the areas of overlap used in this work samples 88 arcmin2, which is about half of the total HST coverage of this field (194 arcmin2). Therefore, we can expect about 2 as many transients if this field were observed again. Full overlapping coverage of this field could yield at least six new Type Ia supernovae, and coupled with spectra and rapid follow-up, can provide essential constraints on fundamental cosmological parameters like the Hubble constant, the mass density of the universe, the cosmological constant, the deceleration parameter, and the age of the Universe (e.g., Riess et al., 1998, 2023). In addition to Type Ia SNe, comprehensive HST coverage of this field could yield at least six CC SNe observations. Particularly when combined with ancillary data, this could contribute significantly to resolving the observed discrepancies in CC SN rates compared to theoretical models (Cappellaro, 2014). With a sufficiently large sample, the SN rate as a function of redshift could also be constrained. Furthermore, the increased volume of observations enhances the likelihood of capturing rare or unexpected events, potentially leading to groundbreaking discoveries in the field.
@@
At least three of the transients are unlikely to be SNe. They either were detected in both epochs or they are isolated with no apparent host galaxy. When considering only the optical emission, transients T7 and T10 exhibit substantial changes in brightness between the two epochs, of = 0.83 0.06 mag and 0.82 0.09 mag, respectively. That would be equivalent to the presence of an additional source with 24.82 and 25.52 mag. The observed centroid positions in both epochs match to within 0001, such that the additional signal is indistinguishable from being due to a nuclear source that changed brightness. We therefore suggest that T7 and T10 may be quasars. The fact that they both were detected in soft X-rays at flux levels 210-15 erg s-1 cm-2 lends additional evidence to this interpretation.
@@
T11 is a unique transient with no visible host, yet was detected in hard X-rays by NuSTAR. If that X-ray emission is indeed associated with T11 then it could be a faint ( 29 mag), high-redshift quasar that briefly flared up to be detectable in F606W. If the X-ray emission is unrelated, then we speculate that T11 may represent a SN located within an undetected host galaxy. That host could be a faint dwarf galaxy at 6, or a more massive host at 6 to explain the non-detection in F606W in the second visit. Unfortunately, T11 falls outside the area with JWST coverage, so we cannot distinguish between these two scenarios at this time. Chakrabarti et al. (2018) have demonstrated that the rates of SNe in dwarf galaxies can be three times higher (per unit volume) than in typical spiral galaxies. Given the limited understanding of star formation in dwarf galaxies, a substantial number of SN candidates like T11 could offer a unique opportunity to identify otherwise undetected dwarf galaxies and gain more insights into their star formation rates.
@@ 5.2 Properties of Variable Candidates
The observed stellar mass distribution (Figure 8b) of galaxies exhibiting variability shows an apparent excess at both 108 and 1010 . We do not detect variability in the most massive galaxies, either due to small number statistics (most likely) or because their SMBHs are less likely to accrete at a sufficient rate to be detectable or have a longer period of relative quiescence between major accretion events. We also do not detect variability in galaxies with 107 , likely due to a combination of both small number statistics and larger photometric uncertainties in these faint systems, such that our 3 threshold is less likely to be met for a given amplitude of variability. If the relative distribution is taken at face values, then it could also hint at the presence of two separate populations of sources with variability: one associated with lower mass 108 hosts, and another associated with higher mass hosts 1010 . Such could be the case if variability associated with CC SNe favors lower mass host galaxies, and variability associated with SN Ia and AGN favors higher mass hosts.
In Figure 8c and d we see that a significant fraction of the galaxies with variability are best fit with stellar population models characterized by active star formation (young ages 30 Myr and high sSFR 30 Gyr-1). Conversely, there is also a sizable fraction that is quiescent (sSFR 0.01 Gyr-1) with stellar populations older than 1 Gyr. A possible excess appears for intermediate population ages (60–600 Myr).
@@ 5.3 Constraints on SMBH Mass, Accretion and Radiation Lifetimes
The sparse and random time sampling of this field does not allow us to trace the evolution of faint AGN, but may still teach us various properties of them. AGN brightness varies on timescales corresponding to the light crossing time in the accretion disks and gas clouds surrounding SMBHs. For example, for a SMBH with a mass 108 , the timescale for variability cannot be shorter than 1 week in the rest-frame (e.g., Xie et al., 2005). For the majority of sources where the variability is due to AGN activity, that variability originates close to the central SMBH. If the timescale we sample is the minimum timescale of variation, we can provide rough constraints on the SMBH masses following Xie et al. (2005). According to Schwarzschild Black Hole Theory (SBHT), the mass of a black hole ( in ) is equal to the minimum timescale of variation ( in yr) as = 4.291011 . Time intervals between observations in various areas of overlap in the TREASUREHUNT data range from 1 day (0.0027 yr) to 4.78 yr. At the median redshift of our variable candidate sample, , the range in rest-frame timescales becomes 0.001–2.08 yr, giving SMBH masses of 5.1108 8.91011 . SMBH masses 1011 are unreasonable: objects within the region of overlap with a 4.78 yr interval between observations likely vary on much shorter timescales and we did not sample the minimum timescale of variation. Even a more typical time interval of 1 yr between epochs will yield a 1011 SMBH mass. Taking the full range in redshifts sampled into account, we will therefore only claim that in select areas of overlap we can probe SMBHs with masses of (3–9)108 , but with no firm upper limit.
@@
We can also frame our finding that a maximum of 0.42% of the general field galaxy population shows significant variability in the context of the accretion and radiation lifetimes of AGN. If we assume that: 1) all galaxies seen in this field with HST have a central SMBH (e.g., Kormendy & Ho, 2013), 2) the central SMBH is always visible when accreting, and 3) that a currently accreting AGN will be variable, then we would deduce that the AGN in our field are actively accreting for 0.42% of the time. Assumptions 2) and 3) are simplified, as some AGN will be hidden at rest-frame UV–Visible wavelengths by surrounding dust. This fraction could be as large as 2/3 of all AGN, if we use the 30% average Lyman continuum escape fraction of weak AGN (Smith et al., 2018, 2020, 2024) as proxy for the fraction of AGN with direct unobscured sight lines to the observer. Figure 8c implies an average SED age of 108 years (100 Myr) for the parent population, although the spread in age is wide (0.01–10 Gyr). A 0.42% fraction of visibly variable AGN would correspond to an implied average AGN activity lifetime of 0.42%108 4105 yr (and a range of 4104–4107 yr). For comparison, Rawes et al. (2015) estimated the optical synchrotron electron life times of AGN with visible jets observed with HST and Chandra at 104 yr. However, these authors state that their estimated synchrotron lifetimes may be too short by at least a factor of two, and may be longer if synchrotron electrons are re-accelerated in the ambient magnetic field. On the other end of the activity scale, Jakobsen et al. (2003) show that at least some quasars can remain active more or less continuously for 10 Myr. With these assumptions and significant uncertainties, a visible AGN accretion time of 104 to 107 yr could indeed result in a 0.42% variability fraction in a galaxy population characterized by an average SED age of 100 Myr (and a range of 0.01–10 Gyr) whose stars and gas feed that central engine.
@@ 5.4 Caveats and Estimated Population Size
We also note that the variable candidates presented here are a strict lower bound to the number present in the field: many genuine variable sources, even if at 6 and detectable in F606W, were certainly missed because they did not vary sufficiently between specific individual visits to meet our 3 threshold, or if the timescale of their variability exceeded the duration of the TREASUREHUNT program. In the present study, the photometric uncertainties in each epoch of observation contributed equally to that 3 threshold. Any future additional observations to similar depths would reduce the photometric uncertainty in the reference magnitudes, thus lowering the variability detection threshold. The magnitude and level of variability for all sources in this work is collected in Table 5, so that they can be compared in future studies within the JWST NEP TDF.